The mgf of a random variable X has the following form: e-8t et 5 Mx(t) = 0.64 . Find ElYX). Answer:-0.2
Answer #1

as we know that mgf of a discrete varible Mx(t) =\sumP(x=x)*etx

comparing it with given MGF:

P(X=0) =0.6

P(X=-8) =1/5 =0.2

P(X=1) =1/5 =0.2

hence E(\sqrt[3]{X} ) =\sumxP(x) =(-8)1/3*0.2+01/3*0.6+(1)1/3*0.2 =-0.4+0.2 =-0.2